OpenStudy (anonymous):

A ball is dropped from the top of a 1,000 foot building. The height of the ball is half its original height after each bounce. What will the height of the ball be after 12 bounces? Using complete sentences, explain the procedure taken to answer this question. (8 points) The answer is 0.48828 but i need to knwo how they got it.

OpenStudy (apoorvk):

This is like a GP isn't it? initial height = a = 1000 the height gets halved after every bounce, so common ratio = 1/2 no. of bounces after which we need to know the height = n = 12 So, find out the 12th term of the GP with first term 100 and common ratio 1/2

OpenStudy (anonymous):

@countrygal is that really the answer???

OpenStudy (anonymous):

x=(1000)(0.5)^12-1 0.48828

OpenStudy (apoorvk):

that answer is right what they posted.

OpenStudy (anonymous):

x=ar^n-1 this is the formula

OpenStudy (anonymous):

ok i got it thasnk yall